# L’approximation arithmétique de Fujita sur courbes adéliques

### v2

@@
Abstract
@@ 1 Introduction
We say that $L$ is a big line bundle if $\operatorname{vol}(L)>0$. Fujita’s approximation theorem states that, for all $\epsilon>0$, there exist a $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(L^{\otimes p})$ can be decomposed as the tensor product of an ample line bundle $A$ and an effective line bundle $M$, which satisfy
@@
This approximated decomposition theorem is proved by T. Fujita [12] for the case of characteristic 00, and by S. Takagi [24] for an arbitrary characteristic. We refer the readers to [17, §11.4] for a survey on this topic.
@@ 1.1 Arithmetic analogue over number fields
Let $K$ be a number field, $\mathscr{X}$ be an arithmetic variety over $\operatorname{Spec}\mathcal{O}_{K}$ of dimension $d$, and $\overline{\mathscr{L}}$ be a Hermitian line bundle on $\mathscr{X}$. We define
@@
$\widehat{\operatorname{vol}}(\overline{\mathscr{L}})=\limsup_{n\to\infty}\frac{\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes n})}{n^{\dim(\mathscr{X})}/\dim(\mathscr{X})!}.$
By the works mentioned above, the arithmetic volume function $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ has many similar properties to those of the geometric setting $\operatorname{vol}(\raisebox{1.72218pt}{.})$, so it is expected that an arithmetically big line bundle $\overline{\mathscr{L}}$, which means $\widehat{\operatorname{vol}}(\overline{\mathscr{L}})>0$, also has the property of Fujita approximation. In fact, this used to be a conjecture of A. Moriwaki.
@@ 1.2 Arakelov geometry over adelic curves
Under this setting, we can generalize many classic results and subjects of Arakelov geometry, including the arithmetic volume function, arithmetic positivity, and Hilbert-Samuel formula. Since the adelic structure becomes more complicated than those of classic cases, the obstructions original from measurability, integrability, and non-discreteness make the developing process more technique. In fact, some classic properties are not valid any longer in this framework, though we have developed many in [9, 10, 11], and see also [23].
@@
Inspired by [4, 26], it is quite natural to ask whether the Fujita approximation is still valid over adelic curves. Over a general adelic curve, we are lack of the Minkowskian property, so we cannot apply the filtrated method by successive minina directly in [4] At the same time, some constructions do not work any longer without the help of integral models.
In [11, §8.5], H. Chen and A. Moriwaki proved a similar result to the arithmetic Fujita approximation over adelic curves, which is called the relative Fujita approximation. More precisely, let $\overline{L}$ be an adelic line bundle, where $L$ is big on $X$. Then for any $t<\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$, there exists a sufficiently large $p\in\mathbb{N}$ and a birational morphism $\varphi:X^{\prime}\rightarrow X$, such that $(\varphi^{*}\overline{L})^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{N}$ is relatively ample and $\overline{M}$ is effective, and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})>pt$. Here $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ denote the asymptotic maximal and minimal slopes respectively.
@@ 1.3 Main result
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, which satisfies either $\mathcal{A}$ is discrete or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be an arithmetically big $\mathbb{R}$-line bundle on $X$ over $S$. Then for any $\epsilon>0$, there exist a sufficiently large $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(\overline{L}^{\otimes p})=\overline{A}\otimes\overline{M}$, where $\overline{A}$ is arithmetically ample and $\overline{M}$ is an arithmetically effective $\mathbb{R}$-line bundle. Moreover, we have
@@
In order to study some arithmetic positivities by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$, we need to suppose the base field $K$ is perfect. For example, for an adelic line bundle $\overline{L}$ where $L$ is nef, we need to apply $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})\geqslant 0$ to deduce $\overline{L}$ is arithmetically nef by [11, Proposition 9.1.7], where we need the perfect condition. See [11, §9] for more details of the necessity.
Of course, in [4, 26], there is no such problem since for the case of number field, because $\mathcal{A}$ is discrete in the language of adelic curve and the base field is perfect by its characteristic 00.
@@ Acknowledgements
I would like to thank Prof. Atsushi Moriwaki for introducing this subject to me. In addition, I would like to thank Prof. Huayi Chen for many useful discussions and suggestions, and Dr. Wenbin Luo for the advices on some key techniques. Part of this work was accomplished during my visit to Institute for Theoretical Sciences, Westlake University, and I would like to thank their generous invitation and reception. Chunhui Liu was supported by National Science Foundation of China No. 12201153.
@@ 2.1.1 Adelic curve
Let $K$ be an arbitrary field, and |.$|\raisebox{1.72218pt}{.}|_{0}$ be the trivial absolute value over $K$, which means
$|a|_{0}=\begin{cases}1,\text{ if }a\neq 0,\\ 0,\text{ if }a=0.\end{cases}$
We note $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ the adelic curve which counts the absolute value |.$|\raisebox{1.72218pt}{.}|_{0}$ once. By definition, $S_{0}$ is a proper adelic curve.
Example 2.1.1 will be used in the article later. We refer the readers to [9, §3.2] for more examples of adelic curve.
@@ 2.1.2 Adelic vector bundle
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, and $E$ be a vector space over $K$. For a family of norms (.$(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ on $E_{w}=E\otimes_{K}K_{w}$ over $K_{w}$, if the norm family (.$(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ on $E$ and its dual norm family (.$(\|\raisebox{1.72218pt}{.}\|^{\vee}_{w})_{w\in\Omega}$ on $E^{\vee}$ are both $\mathcal{A}$-measurable on $\Omega$, then we say $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ is an adelic vector bundle over $S$. If $\dim(E)=1$, we say that $\overline{E}$ is an adelic line bundle over $S$.
We say that an adelic vector bundle $\overline{F}=(F,(\|\raisebox{1.72218pt}{.}\|^{\prime}_{w})_{w\in\Omega})$ is an adelic sub-bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $F$ is a $K$-vector subspace of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime}_{w}$ is the restriction of .$\|\raisebox{1.72218pt}{.}\|_{w}$ on $F$ for each $w\in\Omega$.
An adelic bundle $\overline{G}=(G,(\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w})_{w\in\Omega})$ is said to be an adelic quotient bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $G$ is a quotient vector space of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w}$ is the quotient norm .$\|\raisebox{1.72218pt}{.}\|_{w}$ with respect to the canonical quotient homomorphism $E\rightarrow G$ for each $w\in\Omega$. If $G$ is the quotient space of $E$ modulo $F$, and is equipped with the quotient metrics from $\overline{F}$, then we denote $\overline{G}=\overline{E}/\overline{F}=\overline{E/F}$.
For an adelic vector bundle $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, we define $\det(\overline{E})=(\det(E),(\|\raisebox{1.72218pt}{.}\|_{\det,w})_{w\in\Omega})$ as follows: the $K$-vector space $\det(E)=\bigwedge^{\dim(E)}(E)$ is the determinant space of $E$, and .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$ is the determinant norm on $\det(E)_{w}$ of .$\|\raisebox{1.72218pt}{.}\|_{w}$ for each $w\in\Omega$. We refer the readers to [9, §1.2.3] for the details of the construction of .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$.
@@ 2.2.2 A measure induced by $\mathbb{R}$-filtration
where $\delta_{\lambda_{i}}$ is the Dirac operator at the real number $\lambda_{i}$.
@@ 2.2.3 Harder-Narasimhan filtration
The last equality in Definition 2.2.3 if from [9, Proposition 4.3.36]. In fact, the Harder-Narasimhan filtration of $\overline{E}$ corresponds to an increasing flag of vector subspaces
@@ 2.3 Filtrations on graded algebras
In this part, we study some asymptotic properties of graded linear algebras by $\mathbb{R}$-filtraion.
@@ 2.3.1 Approximable graded algebras
    where $S^{n}B_{p}\rightarrow E_{np}$ is the canonical homomorphism defined by the algebra structure on $E_{\bullet}$.
@@ 2.3.2 Convergence of measures on an approximable algebra
where the function $\lambda(\raisebox{1.72218pt}{.})$ is defined at (1) and the homogeneous degrees of $x_{1}$ and $x_{2}$ are ignored, then we say $E_{\bullet}$ is filtrated with respect to $\mathcal{F}$.
@@
With all the above notations. Suppose that $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ is a filtrated approximable graded algebra with respect to $\lambda(\raisebox{1.72218pt}{.})$ defined at (1), and
@@
where $\lambda_{\max}(\raisebox{1.72218pt}{.})$ is defined at (2). Then the sequence $\left(\frac{\lambda_{\max}(E_{n})}{n}\right)_{n\in\mathbb{N}^{+}}$ converges in $\mathbb{R}$, and the sequence of measures $\left(T_{\frac{1}{n}}(\nu_{E_{n}})\right)_{n\geqslant 1}$ converges weakly to a Borel probability measure on $\mathbb{R}$.
@@ 2.3.3 Limits of a measure from filtrated algebras
where $\lambda_{+}(\raisebox{1.72218pt}{.})$ is defined at (2). See [4, §2.4] for more details.
@@ 3.1.1 Adelic line bundle
|.$\begin{array}[]{rrcl}|\raisebox{1.72218pt}{.}|_{\varphi_{w}}(x):&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&|s|_{\varphi_{w}}(x)\end{array}$
@@
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}}:&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&\sup\limits_{x\in X^{\operatorname{an}}_{w}}\{|s|_{\varphi_{w}}(x)\}.\end{array}$ (5)
@@ 3.1.2 Semi-positive metric
we say that the metric family $\varphi$ is semi-positive.
@@ 3.1.5 Effectivity
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. If $s\in H^{0}(X,L)$ is a non-zero global section which satisfies $\|s\|_{\varphi_{w}}\leqslant 1$ for all $w\in\Omega$, we say that the global section $s$ is effective.
We say that an adelic line bundle $\overline{L}$ is effective or arithmetically effective, if it admits at least one effective global section by the above definition.
@@ 3.3 Arithmetic invariants of graded linear series
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over a proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ contains a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. In §3.1.1, we defined a norm family on $H^{0}(X,L^{\otimes D})$ by (5) indexed by $\Omega$, which is denoted by $D\varphi$ with no ambiguity. By [9, §6.4.1], .$\|\raisebox{1.72218pt}{.}\|_{D\varphi_{w}}$ is ultrametric for every $w\in\Omega_{f}$. Then $(H^{0}(X,L^{\otimes D}),D\varphi)$ is an adelic vector bundle over $S$.
For every $D\in\mathbb{N}$, let $\overline{E}_{D}$ be an adelic vector sub-bundle $H^{0}(X,L^{\otimes D})$ over $S$. In this case, we say that $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ is a normed graded linear series over $S$ if $E_{D}\cdot E_{D^{\prime}}\subseteq E_{D+D^{\prime}}$ for $D,D^{\prime}\in\mathbb{N}$.
@@
In particular, if $\overline{E}_{D}=\left(H^{0}(X,L^{\otimes D}),D\varphi\right)$ as above, we denote $\widehat{\operatorname{vol}}(\overline{L})=\widehat{\operatorname{vol}}(\overline{E}_{\bullet})$ for simplicity.

### v3

@@
[…]
L’approximation arithmétique de Fujita sur courbes adéliques
Arithmetic Fujita approximation over adelic curves
Chunhui Liu Institute for Advanced Study in Mathematics
Harbin Institute of Technology
150001 Harbin
P. R. China
chunhui.liu@hit.edu.cn
(8 janvier 2026)
Résumé
@@ 1 Introduction
We say that $L$ is a big line bundle if $\operatorname{vol}(L)>0$. Fujita’s approximation theorem states that, for all $\epsilon>0$, there exist a $p\in\mathbb{N}^{+}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(L^{\otimes p})$ can be decomposed as the tensor product of an ample line bundle $A$ and an effective line bundle $M$, which satisfy
@@
This approximated decomposition theorem is proved by T. Fujita [12] for the case of characteristic 0, and by S. Takagi [24] for an arbitrary characteristic. We refer the readers to [17, §11.4] for a survey on this topic.
@@ 1.1 Arithmetic analogue over number fields
Let $K$ be a number field, $\mathscr{X}$ be an arithmetic variety over $\operatorname{Spec}\mathcal{O}_{K}$ of dimension $d$, and $\overline{\mathscr{L}}$ be a Hermitian line bundle on $\mathscr{X}$. For a $D\in\mathbb{N}$, we define
@@
$\widehat{\operatorname{vol}}(\overline{\mathscr{L}})=\limsup_{D\to\infty}\frac{\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes D})}{D^{\dim(\mathscr{X})}/\dim(\mathscr{X})!}.$
[…]
@@ 1.2 Arakelov geometry over adelic curves
Under this setting, we are able to generalize many classic results and subjects of Arakelov geometry, including the arithmetic volume function, arithmetic positivity, and Hilbert-Samuel formula. Since the adelic structure becomes more complicated than those of classic cases, the obstructions original from measurability, integrability, and non-discreteness make the developing process more technique. In fact, some classic properties are not valid any longer in this framework, though many of these properties have been developed in [9, 10, 11], and see also [23].
@@
Inspired by [4, 26], it is quite natural to ask whether the Fujita approximation is still valid over adelic curves. Over a general adelic curve, we are lack of the Minkowskian property, so we cannot apply the filtrated method by successive minina directly in [4]. At the same time, some constructions and properties do not work any longer without the help of integral models or Minkowskian property, and see Example 4.1.3 for such an example.
[…]
@@ 1.3 Main result
In the statement of the main result, instead of considering an effective line bundle under the arithmetic setting, we use the notation that a line bundle has a non-zero small section over an adelic curve, and see §3.1.5 for the precise definition.
@@
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, which satisfies either $\mathcal{A}$ is discrete or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be an arithmetically big $\mathbb{R}$-line bundle on $X$ over $S$. Then for any $\epsilon>0$, there exist a sufficiently large $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(\overline{L}^{\otimes p})=\overline{A}\otimes\overline{M}$, where $\overline{A}$ is arithmetically ample and $\overline{M}$ is an adelic $\mathbb{R}$-line bundle have a non-zero small section. Moreover, we have
@@
[…]
Of course, in [4, 26], there is no such problem since for the case of number field, because $\mathcal{A}$ is discrete in the language of adelic curve and the base field is perfect by its characteristic 0.
@@ Acknowledgements
I would like to thank Prof. Atsushi Moriwaki for introducing this subject to me, and suggesting me to use the notation "having a non-zero small section". In addition, I would like to thank Prof. Huayi Chen for many useful discussions and suggestions, and Dr. Wenbin Luo for the advices on some key techniques. In particular, I would like to thank the anonymous referee for his or her valuable comments on this article. Part of this work was accomplished during my visit to Institute for Theoretical Sciences, Westlake University, and I would like to thank their generous invitation and reception. Chunhui Liu was supported by National Science Foundation of China No. 12201153.
@@ 2.1.1 Adelic curve
Let $K$ be an arbitrary field, and $|\raisebox{1.72218pt}{.}|_{0}$ be the trivial absolute value over $K$, which means
[…]
We note $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ the adelic curve which counts the absolute value $|\raisebox{1.72218pt}{.}|_{0}$ once. By definition, $S_{0}$ is a proper adelic curve.
Example 2.1.1 will be used in the article later. We refer the readers to [9, §3.2] for more examples of adelic curves.
@@ 2.1.2 Adelic vector bundle
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, and $E$ be a vector space over $K$. For a family of norms $(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ on $E_{w}=E\otimes_{K}K_{w}$ over $K_{w}$, if the norm family $(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ on $E$ and its dual norm family $(\|\raisebox{1.72218pt}{.}\|^{\vee}_{w})_{w\in\Omega}$ on $E^{\vee}$ are both $\mathcal{A}$-measurable on $\Omega$, and $(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ is dominated, then we say $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ is an adelic vector bundle over $S$. If $\dim(E)=1$, we say that $\overline{E}$ is an adelic line bundle over $S$. See [9, Definition 4.1.28] for more details on this notion.
We say that an adelic vector bundle $\overline{F}=(F,(\|\raisebox{1.72218pt}{.}\|^{\prime}_{w})_{w\in\Omega})$ is an adelic sub-bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $F$ is a $K$-vector subspace of $E$, and $\|\raisebox{1.72218pt}{.}\|^{\prime}_{w}$ is the restriction of $\|\raisebox{1.72218pt}{.}\|_{w}$ on $F$ for each $w\in\Omega$.
An adelic bundle $\overline{G}=(G,(\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w})_{w\in\Omega})$ is said to be an adelic quotient bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $G$ is a quotient vector space of $E$, and $\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w}$ is the quotient norm $\|\raisebox{1.72218pt}{.}\|_{w}$ with respect to the canonical quotient homomorphism $E\rightarrow G$ for each $w\in\Omega$. If $G$ is the quotient space of $E$ modulo $F$, and is equipped with the quotient metrics from $\overline{F}$, then we denote $\overline{G}=\overline{E}/\overline{F}=\overline{E/F}$.
For an adelic vector bundle $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, we define $\det(\overline{E})=(\det(E),(\|\raisebox{1.72218pt}{.}\|_{\det,w})_{w\in\Omega})$ as follows : the $K$-vector space $\det(E)=\bigwedge^{\dim(E)}(E)$ is the determinant space of $E$, and $\|\raisebox{1.72218pt}{.}\|_{\det,w}$ is the determinant norm on $\det(E)_{w}$ of $\|\raisebox{1.72218pt}{.}\|_{w}$ for each $w\in\Omega$. We refer the readers to [9, §1.1.13, §1.2.3] for the details of the construction of $\|\raisebox{1.72218pt}{.}\|_{\det,w}$.
@@ 2.2.2 A measure induced by $\mathbb{R}$-filtration
where $\delta_{\lambda_{i}}$ is the Dirac measure at $\lambda_{i}\in\mathbb{R}$.
@@ 2.2.3 Harder-Narasimhan filtration
The last equality in Definition 2.2.3 is from [9, Proposition 4.3.36]. In fact, the Harder-Narasimhan filtration of $\overline{E}$ corresponds to an increasing flag of vector subspaces
@@ 2.3 Filtrations on graded algebras
In this part, we study some asymptotic properties of graded linear algebras by $\mathbb{R}$-filtration.
@@ 2.3.1 Approximable graded algebras
    where $S^{n}E_{p}\rightarrow E_{np}$ is the canonical homomorphism defined by the algebra structure on $E_{\bullet}$.
@@ 2.3.2 Convergence of measures on an approximable algebra
[…]
@@
[…]
@@
[…]
@@ 2.3.3 Limits of a measure from filtrated algebras
[…]
@@ 3.1.1 Adelic line bundle
$\begin{array}[]{rrcl}|\raisebox{1.72218pt}{.}|_{\varphi_{w}}(x):&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&|s|_{\varphi_{w}}(x)\end{array}$
@@
$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}}:&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&\sup\limits_{x\in X^{\operatorname{an}}_{w}}\{|s|_{\varphi_{w}}(x)\}.\end{array}$ (5)
@@ 3.1.2 Semi-positive metric
we say that the metric family $\varphi$ is semi-positive. See [9, Definition 2.2.21] for the details of the construction of a quotient norm family.
@@ 3.1.5 Effectivity
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. If $s\in H^{0}(X,L)$ is a non-zero global section which satisfies $\|s\|_{\varphi_{w}}\leqslant 1$ for all $w\in\Omega$, we say that the global section $s$ is small, or $s$ is a small section.
We say that an adelic line bundle $\overline{L}$ has a non-zero small section, if it admits at least one small global section by the above definition.
[…]
In [11, Definition 8.5.4], the above $\overline{L}$ is said to be effective in the arithmetic sense, but this notion is only used in §8.5 and §8.6 of [11]. In order to avoid the confusion, we use the word $\overline{L}$ has a non-zero small section” instead of [11, Definition 8.5.4] in this article, because some multiple of an arithmetically big adelic line bundle is not necessarily effective. This is different from the classic case over a number field.
@@ 3.3 Arithmetic invariants of graded linear series
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over a proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ contains a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. In §3.1.1, we defined a norm family on $H^{0}(X,L^{\otimes D})$ by (5) indexed by $\Omega$, which is denoted by $D\varphi$ with no ambiguity. By [9, §6.4.1], $\|\raisebox{1.72218pt}{.}\|_{D\varphi_{w}}$ is ultrametric for every $w\in\Omega_{f}$. Then $(H^{0}(X,L^{\otimes D}),D\varphi)$ is an adelic vector bundle over $S$.
For every $D\in\mathbb{N}$, let $\overline{E}_{D}$ be an adelic vector sub-bundle of $H^{0}(X,L^{\otimes D})$ over $S$. In this case, we say that $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ is a normed graded linear series over $S$ if $E_{D}\cdot E_{D^{\prime}}\subseteq E_{D+D^{\prime}}$ for $D,D^{\prime}\in\mathbb{N}$.
@@
In particular, if $\overline{E}_{n}=\left(H^{0}(X,L^{\otimes n}),n\varphi\right)$ as above, we denote $\widehat{\operatorname{vol}}(\overline{L})=\widehat{\operatorname{vol}}(\overline{E}_{\bullet})$ for simplicity.